# Identifiability of nonlinear ODE Models with Time-Varying Parameters: the General Analytical Solution and Applications in Viral Dynamics

### v4

@@ 4.4.3 Identifiability
We use (4.6) to compute ${}^{u}\chi$. Note that, in this case $m=m_{w}=1$, and, from (3.8), ${}^{m}\nu^{1}_{1}=\nu^{1}_{1}=\frac{\rho}{\sin(\theta-\phi)}$. We obtain:
@@ 5.1 Indistinguishable states and parameters
Let us consider the system $\Sigma^{\prime}$ returned by Algorithm 1. It still satisfies Equation (2.2). Let us assume that the initial time is $t=0$ and the initial state is $x(0)=x_{0}$. The state at time $t$ is obtained by solving the following differential equation (where all the quantities are the ones that characterize the system $\Sigma^{\prime}$ returned by Algorithm 1):
@@ 5.2 Minimal missing information to achieve observability/identifiability
In particular, we would like an automatic answer, which holds for any system, as in the case of the observability and identifiability analysis that are obtained by simply running Algorithm 1 and 4.
@@ 6.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides:
$\Omega=\textnormal{span}\{\nabla h_{1},~{}\nabla h_{2}\}$
$=\textnormal{span}\{[0,0,1,0,0,0,0,0],~{}[1,1,0,0,0,0,0,0]\}.$
$\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ vanishes). The system is not in canonical form with respect to its unknown input. We enter the loop at lines 9-14.
Line 11 provides no functions ($m=0$). Line 12 provides: ${}^{m}\mu^{0}_{0}=~{}^{m}\nu^{0}_{0}=1$ obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: ${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}=g^{0}$. Line 13 computes
$\Omega=\left<\left.\left\{\begin{array}[]{l}g^{0}\\ g^{1}\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2}\right\}\right>,$
as ${}^{m}\widehat{g}^{0}=g^{0}$. By a direct computation, we obtain:
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=1=m_{w}$. Hence, the second condition at Line 9 is not met and the loop is interrupted. In addition, as the condition at Line 15 is met, also the main loop is interrupted. The execution of Algorithm 1 continues with Line 45. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
(we would obtain the same final result by setting $\widetilde{h}_{1}=h_{3}=\mathcal{L}_{g^{0}}h_{1}$). Then, Line 46 provides:
@@
Finally, from (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 47, namely:
@@ 6.4 Identifiability analysis
In addition, from Equation (6.4) and by knowing that $m=m_{w}=1$ (canonic system), we have ${}^{m}\nu^{1}_{1}=\nu^{1}_{1}=-\frac{1}{T_{U}V(\delta-\rho)}$.
@@ 7.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides $\Omega=\textnormal{span}\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3}\}$, and $\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=\mathcal{L}_{g^{1}}h_{3}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2},~{}h_{3}$ vanishes). The system is not in canonical form with respect to its unknown input. We enter the loop at lines 9-14. Line 11 provides no functions ($m=0$). Line 12 provides: ${}^{m}\mu^{0}_{0}=~{}^{m}\nu^{0}_{0}=1$ obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: ${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}=g^{0}$.
Line 13 computes
$\Omega=\left<\left.\left\{\begin{array}[]{l}g^{0}\\ g^{1}\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3}\right\}\right>,$
as ${}^{m}\widehat{g}^{0}=g^{0}$. By a direct computation, we obtain:
$\Omega=\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3},~{}\nabla\mathcal{L}_{g^{0}}h_{1},~{}\nabla\mathcal{L}_{g^{0}}h_{2}\right\}=$
$\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3},~{}\nabla h_{4},~{}\nabla h_{5}\right\}$
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=1=m_{w}$. Hence, the second condition at Line 9 is not met and the loop is interrupted. In addition, as the condition at Line 15 is met, also the main loop is interrupted. The execution of Algorithm 1 continues with Line 45. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
(we would obtain the same final result by setting $\widetilde{h}_{1}=h_{5}=\mathcal{L}_{g^{0}}h_{2}$). Then, Line 46 provides:
@@
Finally, from (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 47, namely:
@@ 7.4.1 First symmetry
${}^{u}\chi=-~{}^{m}\nu^{1}_{1}(\xi^{0}_{1}+\xi^{1}_{1}w_{1})=-\frac{\xi^{1}_{1}}{S\gamma p(A+I)}w_{1}=-\frac{1}{S}\beta$
because ${}^{m}\nu^{1}_{1}=\nu^{1}_{1}=\frac{1}{S\gamma p(A+I)}$. Hence, we have the following first symmetry of the unknown input:
@@ 7.4.2 Second symmetry
${}^{u}\chi=-~{}^{m}\nu^{1}_{1}(\xi^{0}_{1}+\xi^{1}_{1}w_{1})=-\frac{\xi^{0}_{1}}{S\gamma p(A+I)}-\frac{\xi^{1}_{1}}{S\gamma p(A+I)}w_{1}=$
@@
because ${}^{m}\nu^{1}_{1}=\nu^{1}_{1}=\frac{1}{S\gamma p(A+I)}$. Hence, we have the following second symmetry of the unknown input:
@@ 8.1 Implementation of Algorithm 1
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides:
@@
$\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=\mathcal{L}_{g^{2}}h_{1}=\mathcal{L}_{g^{2}}h_{2}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ vanishes). The system is not in canonical form with respect to its unknown input. We enter the loop at lines 9-14.
Line 11 provides no functions ($m=0$). Line 12 provides: ${}^{m}\mu^{0}_{0}=~{}^{m}\nu^{0}_{0}=1$ obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: ${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}=g^{0}$. Line 13 computes
$\Omega=\left<\left.\left\{\begin{array}[]{ll}g^{0}&\\ g^{1}&g^{2}\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2}\right\}\right>,$
as ${}^{m}\widehat{g}^{0}=g^{0}$. By a direct computation, we obtain:
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=2=m_{w}$. Hence, the second condition at Line 9 is not met and the loop is interrupted. In addition, as the condition at Line 15 is met, also the main loop is interrupted.
The execution of Algorithm 1 continues with Line 45. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
Then, Line 46 provides:
@@
From (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 47, namely:

### v5

@@ 4.4.3 Identifiability
We use (4.7) to compute ${}^{u}\chi$. Note that, in this case $m=m_{w}=1$, and, from (3.12), $\nu^{1}_{1}=\nu^{1}_{1}=\frac{\rho}{\sin(\theta-\phi)}$. We obtain:
@@ 5.1 Indistinguishable states and parameters
Let us consider the system $\mathcal{E}$ returned by Algorithm 1. It still satisfies Equation (2.2). Let us assume that the initial time is $t=0$ and the initial state is $x(0)=x_{0}$. The state at time $t$ is obtained by solving the following differential equation (where all the quantities are the ones that characterize the system $\mathcal{E}$ returned by Algorithm 1):
@@ 5.2 Minimal missing information to achieve observability/identifiability
In particular, we would like an automatic answer, which holds for any system, as in the case of the observability and identifiability analysis that are obtained by simply running Algorithm 1 and 5.
@@ 6.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 1 provides:
$\Omega=\textnormal{span}\{\nabla h_{1},~{}\nabla h_{2}\}=\textnormal{span}\{[0,0,1,0,0,0,0,0],~{}[1,1,0,0,0,0,0,0]\}.$
$\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ vanishes). The system is not in canonical form with respect to its unknown input.
We have $m=0$ and $\mu^{0}_{0}=~{}\nu^{0}_{0}=1$ obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: $\widehat{g}^{0}=~{}\nu^{0}_{0}g^{0}=g^{0}$, and $\widehat{g}^{1}=g^{1}$.
Algorithm 2 provides Finish=False. We have:
[…]
as $\widehat{g}=\widehat{g}^{0}=g^{0}$. By a direct computation, we obtain:
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=1=m_{w}$. Hence, the execution of Algorithm 1 continues with Line 18. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
(we would obtain the same final result by setting $\widetilde{h}_{1}=h_{3}=\mathcal{L}_{g^{0}}h_{1}$). Then, Line 19 provides:
@@
Finally, from (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 20, namely:
@@ 6.4 Identifiability analysis
In addition, from Equation (6.4) and by knowing that $m=m_{w}=1$ (canonic system), we have $\nu^{1}_{1}=\nu^{1}_{1}=-\frac{1}{T_{U}V(\delta-\rho)}$.
@@ 7.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 1 provides $\Omega=\textnormal{span}\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3}\}$, and $\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=\mathcal{L}_{g^{1}}h_{3}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2},~{}h_{3}$ vanishes). The system is not in canonical form with respect to its unknown input. We have $m=0$ and $\mu^{0}_{0}=~{}\nu^{0}_{0}=1$ obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: $\widehat{g}^{0}=~{}\nu^{0}_{0}g^{0}=g^{0}$, and $\widehat{g}^{1}=g^{1}$.
Algorithm 2 provides Finish=False.
We have:
[…]
as $\widehat{g}=\widehat{g}^{0}=g^{0}$. By a direct computation, we obtain:
$\Omega=\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3},~{}\nabla\mathcal{L}_{g^{0}}h_{1},~{}\nabla\mathcal{L}_{g^{0}}h_{2}\right\}=\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla h_{3},~{}\nabla h_{4},~{}\nabla h_{5}\right\}$
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=1=m_{w}$. Hence, the execution of Algorithm 1 continues with Line 18. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
(we would obtain the same final result by setting $\widetilde{h}_{1}=h_{5}=\mathcal{L}_{g^{0}}h_{2}$). Then, Line 19 provides:
@@
Finally, from (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 20, namely:
@@ 7.4.1 First symmetry
${}^{u}\chi=-~{}\nu^{1}_{1}(\xi^{0}_{1}+\xi^{1}_{1}w_{1})=-\frac{\xi^{1}_{1}}{S\gamma p(A+I)}w_{1}=-\frac{1}{S}\beta$
because $\nu^{1}_{1}=\nu^{1}_{1}=\frac{1}{S\gamma p(A+I)}$. Hence, we have the following first symmetry of the unknown input:
@@ 7.4.2 Second symmetry
${}^{u}\chi=-\nu^{1}_{1}(\xi^{0}_{1}+\xi^{1}_{1}w_{1})=-\frac{\xi^{0}_{1}}{S\gamma p(A+I)}-\frac{\xi^{1}_{1}}{S\gamma p(A+I)}w_{1}=$
@@
because $\nu^{1}_{1}=\nu^{1}_{1}=\frac{1}{S\gamma p(A+I)}$. Hence, we have the following second symmetry of the unknown input:
@@ 8.1 Implementation of Algorithm 1
We run Algorithm 1 to obtain the observability codistribution. Line 1 provides:
@@
$\textnormal{deg}_{w}\left({\Omega}\right)=0~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{1}}h_{2}=\mathcal{L}_{g^{2}}h_{1}=\mathcal{L}_{g^{2}}h_{2}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ vanishes). The system is not in canonical form with respect to its unknown input. We have $m=0$ and $\mu^{0}_{0}=\nu^{0}_{0}=1$, obtained from (3.6) (these tensors have a single entry as the indices only take the value 0). In addition, from (3.7) we obtain: $\widehat{g}^{0}=\nu^{0}_{0}g^{0}=g^{0}$, $\widehat{g}^{1}=g^{1}$, and $\widehat{g}^{2}=g^{2}$.
Algorithm 2 provides Finish=False. In addition, from (3.8), we obtain: $\widehat{g}=\widehat{g}^{0}=g^{0}$. We have:
[…]
By a direct computation, we obtain:
@@
Now, $\textnormal{deg}_{w}\left(\Omega\right)=2=m_{w}$. Hence, the execution of Algorithm 1 continues with Line 18. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
@@
Then, Line 19 provides:
@@
From (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 20, namely: